OpenStudy (anonymous):

An arrow is shot straight up in the air with an initial speed of 240 ft/s. If on striking the ground it embeds itself 8.00 in into the ground, find the magnitude of the acceleration (assumed constant) required to stop the arrow, in units of feet/second^2. not sure where to start on this one

OpenStudy (danjs):

First find the velocity of the arrow when it strikes the ground. The arrow is fired 240 ft/s straight up, reaches a maximum height, then falls straight down. The final velocity when it reaches the ground is then used to find the acceleration needed to go from that velocity to zero, over a distance of 8.00 inches, or 8/12 ft.

OpenStudy (shamim):

We need to use the equation is v^2=u^2+2*a*s Here given v=0 u=240ft/s s=8/12 ft

OpenStudy (shamim):

Feel free to ask for more explanation